A. Find the length of the following confidence interval.

1. Upper limit = 0.995

Lower Limit = 0.437


2. Upper limit = 394.14

Lower Limit = 354.74


3. Upper limit = 0.02946

Lower Limit = 0.02244


4. 0.475 < p < 0.735


5. 0.355 < p < 0.570


1
Expert's answer
2022-05-09T16:08:04-0400

1.


"Length=0.995-0.437=0.558"

2.


"Length=394.14-354.74=39.40"

3.


"Length=0.02946-0.02244=0.00702"

4.


"Length=0.735-0.475=0.260"

5.


"Length=0.570-0.355=0.215"